# Monte Carlo Tree Diffusion for System 2 Planning

### v2

@@ 1 Introduction
Diffusion models have recently emerged as a powerful approach to planning, enabling the generation of complex trajectories by modeling trajectory distributions using large-scale offline data (Janner et al., 2022; Ajay et al., 2022; Zhou et al., 2024; Chen et al., 2024a, c, b). Unlike traditional autoregressive planning methods, diffusion-based planners, such as Diffuser (Janner et al., 2022), generate entire trajectories holistically through a series of denoising steps, eliminating the need for a forward dynamics model. This approach effectively addresses key limitations of forward models, such as poor long-term dependency modeling and error accumulation (Hafner et al., 2022; Hamed et al., 2024), making it particularly well-suited for planning tasks with long horizons or sparse rewards.
Despite their strengths, it remains uncertain how diffusion-based planners can effectively enhance planning accuracy by utilizing additional test-time compute (TTC)—a property referred to as TTC scalability. One potential approach is to increase the number of denoising steps or, alternatively, draw additional samples (Zhou et al., 2024). However, it is known that performance gains from increasing denoising steps plateau quickly (Karras et al., 2022; Song et al., 2020a, b), and independent random searches with multiple samples are highly inefficient as they fail to leverage information from other samples. Moreover, how to effectively manage the exploration-exploitation tradeoff within this framework also remains unclear.
In contrast, Monte Carlo Tree Search (MCTS) (Coulom, 2006), a widely adopted planning method, demonstrates robust TTC scalability. By leveraging iterative simulations, MCTS refines decisions and adapts based on exploratory feedback, making it highly effective in improving planning accuracy as more computation is allocated. This capability has established MCTS as a cornerstone in many System 2 reasoning tasks, such as mathematical problem-solving (Guan et al., 2025; Zhang et al., 2024a) and program synthesis (Brandfonbrener et al., 2024). However, unlike diffusion-based planners, traditional MCTS relies on a forward model for tree rollouts, inheriting its limitations including losing global consistency. In addition to being restricted to discrete action spaces, the resulting search tree can grow excessively large in both depth and width. This leads to significant computational demands, particularly in scenarios involving long horizons and large action spaces.
@@ 2.1 Diffuser: Diffusion Models for Planning
Diffuser (Janner et al., 2022) addresses long‐horizon decision-making by treating entire trajectories as a matrix
@@
Accordingly, at each denoising step, gradient information from nudges the model toward trajectories that appear both feasible (as learned from the offline data) and promising with respect to returns, in a manner akin to classifier guidance in image diffusion (Dhariwal & Nichol, 2021).
Diffusion Forcing (Chen et al., 2024a) extends the above framework by allowing tokenization of . This tokenization allows each token to be denoised at a different noise level, enabling partial denoising of segments where uncertainty is high (e.g., future plans), without requiring a complete transition from full noise to no noise across the entire trajectory. Such token-level control is particularly beneficial in domains that demand causal consistency, such as long-horizon planning problems.
@@ 2.2 Monte Carlo Tree Search
Monte Carlo Tree Search (MCTS) is a planning algorithm that combines tree search with stochastic simulations to effectively balance exploration and exploitation (Coulom, 2006). Typically, MCTS proceeds in four stages: selection, expansion, simulation, and backpropagation. During selection, the algorithm starts at a root node corresponding to the current state (or partial plan) and descends the tree according to a policy such as Upper Confidence Bounds for Trees (UCT) (Kocsis & Szepesvári, 2006), aiming to balancely choose the most promising child at each step or explore unexperienced states. Once a leaf or expandable node is reached, expansion adds new child nodes representing previously unexplored actions or plans.
@@ 3.1 Denoising as Tree-Rollout
To address this issue, we first introduce the Denoising as Tree-Rollout process by leveraging the semi-autoregressive denoising process (Chen et al., 2024a). Specifically, we partition the full trajectory into subplans, such that . Unlike the standard diffuser, where all subplans share the same global denoising schedule, our approach allows assigning independent denoising schedules to each subplan. By applying faster denoising to earlier subplans and slower denoising to later subplans, the process effectively becomes causal and semi-autoregressive. This allows the denoising process to determine the future conditioned on the already determined past. Consequently, while preserving the globally consistent and holistic generation advantages of Diffuser, the denoising process approximates:
Notably, although this formulation appears autoregressive, it is still executed as a single denoising process by controlling noise levels across subplans (Chen et al., 2024a).
@@ 3.2 Guidance Levels as Meta-Actions
For simplicity, consider two guidance levels: guide and no_guide. In MCTD, we observe that sampling from the prior distribution , i.e., using the standard diffusion sampler represents exploratory behavior as it does not attempt to achieve any goal but only imitates the prior behavior contained in the offline data. Thus, we correspond it to the no_guide meta-action. Conversely, sampling from a goal-directed distribution , such as through classifier-guided diffusion (Dhariwal & Nichol, 2021), represents exploitative behavior, and thus meta-action, guide, is assigned. This steers the sampling process toward achieving a specific goal defined by a reward function .
@@ 3.3 Jumpy Denoising as Fast Simulation
In MCTD, we implement this simulation functionality using the fast jumpy denoising process based on DDIM (Song et al., 2020a). Specifically, when the tree-rollout denoising process has progressed up to the -th subplan, the remaining steps are denoised quickly by skipping every step:
@@ 4 Related Works
Diffusion models (Sohl-Dickstein et al., 2015) have recently shown significant promise for long‐horizon trajectory planning, particularly in settings with sparse rewards, by learning to generate entire sequences rather than stepwise actions (Janner et al., 2022; Chen et al., 2024a; Ajay et al., 2022; Zhou et al., 2024; Chen et al., 2024c, b, ). Various enhancements have been proposed to extend their capabilities. For instance, Chen et al.(2024a) incorporate causal noise scheduling for semi‐autoregressive planning, and other works introduce hierarchical structures (Chen et al., 2024c, b, ), low‐level value learning policies (Chen et al., 2024b), or classifier‐free guidance (Ajay et al., 2022; Zhou et al., 2024). Despite these developments, the explicit interplay between exploration and exploitation in diffusion sampling has received relatively little attention.
Chen et al. (2024a) further propose Monte Carlo Guidance (MCG) to leverage multiple sub‐plans and average their guidance signals, thereby biasing denoising toward higher‐reward outcomes. While this can encourage planning toward an expected reward over multiple rollouts, it does not implement an explicit search mechanism. The detailed discussion is in Appendix C. Similarly, Anonymous (2024) apply a discrete diffusion denoising process to tasks such as chess, implicitly modeling MCTS without explicit tree expansion.
MCTS (Coulom, 2006) has achieved impressive results across various decision-making problems, particularly when combined with learned policies or value networks (Silver et al., 2016; Schrittwieser et al., 2020). It has also been applied to System 2 reasoning in Large Language Models (LLMs) to enhance structured problem-solving (Xiang et al., 2025; Yao et al., 2024; Zhang et al., 2024b). However, to the best of our knowledge, this work is the first to integrate tree search with diffusion models for full trajectory generation, bridging structured search with generative planning.
@@ 5 Experiments
We evaluate the proposed approach, MCTD, on a suite of tasks from the Offline Goal‐conditioned RL Benchmark (OGBench) (Park et al., 2024), which spans diverse domains such as maze navigation with multiple robot morphologies (e.g., point‐mass or ant) and robot‐arm manipulation. Our chosen tasks—point and antmaze navigation, multi‐cube manipulation, and a newly introduced visual pointmaze—jointly assess a planner’s ability to handle long‐horizon planning, sequential manipulation, and partial visual observability. In the visual pointmaze, an agent perceives RGB image observations of the 3D environment, thereby testing each method’s resilience to partial observability and the ability to handle image-based planning.
@@ 5.1 Baselines
We compare MCTD against several baselines, each employing a distinct strategy for test‐time computation. First, we include the standard single‐shot Diffuser (Janner et al., 2022), which plans only at the start of an episode and executes that plan without further adjustments. Next, we consider two methods that allocate additional test‐time compute differently. Diffuser‐Replanning periodically replans at fixed intervals, allowing partial corrections during the episode. This enables us to gauge the benefits of iterative replanning without a full tree search. Diffuser‐Random Search (Zhou et al., 2024) generates multiple random trajectories in parallel and selects the highest‐scoring one according to the same reward used by MCTD. While this increases test‐time sampling and applies multiple guidance levels for different samples, it lacks the systematic expansion and pruning of a tree‐search algorithm. We also compare our method to Diffusion Forcing (Chen et al., 2024a), which introduces a causal denoising schedule that enables semi‐autoregressive trajectory generation. This comparison helps isolate the benefits of semi‐autoregressive denoising from our full tree‐structured approach. Further implementation details about all baselines can be found in Appendix A.1.
@@ 5.2 Maze Navigation with Point‐Mass and Ant Robots
We begin with the pointmaze and antmaze tasks from OGBench, featuring mazes of varying sizes (medium, large, and giant). The agent is rewarded for reaching a designated goal region, and the relevant dataset (navigate) comprises long‐horizon trajectories that challenge each method’s capacity for exploration. As in prior work (Janner et al., 2022), we employ a heuristic controller for the pointmaze environment, whereas the antmaze environment uses a value‐learning policy (Chen et al., 2024b) for low‐level control. Appendix A.5 provides more details on this task evaluation.
@@ Results.
In the antmaze domain, MCTD continues to perform nearly perfectly, while baseline performance generally degrades more significantly than in pointmaze, except for the medium-sized settings under Diffusion Forcing. A principal reason for the baseline’s degradation is that the value learning policy (Wang et al., 2022) struggles in high‐dimensional control tasks. Even the performance degradations exist due to inaccurate execution, MCTD consistently outperforms baselines and the gap is larger as the maze size is larger.
@@ 5.3 Robot Arm Cube Manipulation
For multi‐cube manipulation tasks from OGBench, a robot arm must move one to four cubes to specific table locations as shown in Figure 4a. Increasing the number of cubes grows both the planning horizon and the combinatorial complexity. While the planner (MCTD or a baseline) issues high‐level positions for the cubes, a value-learning policy (Wang et al., 2022) executes local actions. We augment MCTD with object‐wise guidance by selecting which cube is manipulated at each step, avoiding simultaneous movements of multiple cubes. We note that the object-wise guidance is particularly suited to MCTD because the tree-expansion mechanism naturally supports discrete ”meta-actions” at each node, such as selecting which object to move next. Single-pass diffusion methods, by contrast, attempt to produce a single, holistic trajectory in one go, leaving no opportunity to decide object-by-object. This guidance and additional task‐specific reward shaping are described in Appendix A.6.
@@ Results.
Table 2 shows success rates for one to four cubes. All methods perform comparably on single‐cube tasks, but performance drops significantly for multi‐cube scenarios, especially for standard diffusion baselines that struggle to decide cube ordering and handle subtasks such as stacking. However, MCTD-Replanning shows significant performance gaps from other models for multi-cube scenarios. MCTD exhibits moderate gains (e.g., 22% on two cubes), though it suffers from holistic plan entanglements when multiple objects are involved as shown in Figure 4b. To address this, MCTD‐Replanning periodically replans, effectively separating each cube’s movements as shown in Figure 4c. This boosts success on the double‐cube problem from 22% to 50%, and also benefits tasks with more cubes. Another challenge arises from the value-learning policy, which generalizes poorly as the number of cubes increases. Even if MCTD generates a suitable global plan, flawed local control undermines the final outcome. Interestingly, training the policy solely on single‐cube data still allows MCTD‐Replanning to solve a portion (24%) of quadruple‐cube scenarios, illustrating the potential of object‐wise guidance and iterative replanning.
@@ 5.4 Visual Pointmaze
We pre‐encode observations with a Variational Autoencoder (VAE) (Kingma & Welling, 2013) to generate latent states suitable for diffusion‐based planning. Because Diffuser’s default heuristic controller is inapplicable to purely visual inputs, we design an inverse dynamics model that maps consecutive latent observations to actions, . For guidance, we employ a position estimator trained on ground‐truth coordinates; however, this estimator does not update the underlying VAE, ensuring that the planner remains in a partially observable regime. The details for this task setting are discussed in Appendix A.7.
@@ 7 Conclusion
Acknowledgements
This research was supported by GRDC (Global Research Development Center) Cooperative Hub Program (RS-2024-00436165) and Brain Pool Plus Program (No. 2021H1D3A2A03103645) through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT.
@@ A.1 Baselines
Diffuser (Janner et al., 2022): We use the standard single‐shot Diffuser as our primary diffusion baseline. It generates a full trajectory by denoising from maximum to zero noise in one pass, guided by a return‐predicting function . While it can produce coherent plans, it lacks any mechanism to adapt or refine these plans at test time.
@@
Diffuser-Random Search: We consider a random‐search variant of Diffuser that generates multiple trajectories from the randomly noised plans and picks the best candidate according to a heuristic or learned value function. This approach can be seen as a “Sample-Score-Rank” method proposed in (Zhou et al., 2024), increasing test‐time compute by drawing more samples, but lacking the systematic exploration or adaptation found in tree‐search paradigms.
Diffusion Forcing: Diffusion Forcing (Chen et al., 2024a) extends Diffuser by introducing a tokenized, causal noise schedule. Subplans of the trajectory can be denoised at different rates, allowing partial re‐sampling of specific time segments without re‐optimizing the entire plan. In our experiments, we evaluate Diffusion Forcing with the Transformer backbone provided in their official repository, which performed better in our preliminary tests. This method helps isolate the benefit of tokenized partial denoising from our tree‐structured approach. We note that Diffusion Forcing baselines are basically designed to do replanning, so we follow the setting in this evaluation.

### v3

@@ 1 Introduction
Diffusion models have recently emerged as a powerful approach to planning, enabling the generation of complex trajectories by modeling trajectory distributions using large-scale offline data (diffuser; decisiondiffuer; d-mpc; diffusion_forcing; hd; plandq). Unlike traditional autoregressive planning methods, diffusion-based planners, such as Diffuser (diffuser), generate entire trajectories holistically through a series of denoising steps, eliminating the need for a forward dynamics model. This approach effectively addresses key limitations of forward models, such as poor long-term dependency modeling and error accumulation (director; drstrategy), making it particularly well-suited for planning tasks with long horizons or sparse rewards.
Despite their strengths, it remains uncertain how diffusion-based planners can effectively enhance planning accuracy by utilizing additional test-time compute (TTC)—a property referred to as TTC scalability. One potential approach is to increase the number of denoising steps or, alternatively, draw additional samples (d-mpc). However, it is known that performance gains from increasing denoising steps plateau quickly (karras2022elucidating; ddim; song2020score), and independent random searches with multiple samples are highly inefficient as they fail to leverage information from other samples. Moreover, how to effectively manage the exploration-exploitation tradeoff within this framework also remains unclear.
In contrast, Monte Carlo Tree Search (MCTS) (mcts), a widely adopted planning method, demonstrates robust TTC scalability. By leveraging iterative simulations, MCTS refines decisions and adapts based on exploratory feedback, making it highly effective in improving planning accuracy as more computation is allocated. This capability has established MCTS as a cornerstone in many System 2 reasoning tasks, such as mathematical problem-solving (guan2025rstar; zhang2024accessing) and program synthesis (brandfonbrener2024vermcts). However, unlike diffusion-based planners, traditional MCTS relies on a forward model for tree rollouts, inheriting its limitations including losing global consistency. In addition to being restricted to discrete action spaces, the resulting search tree can grow excessively large in both depth and width. This leads to significant computational demands, particularly in scenarios involving long horizons and large action spaces.
@@ 2.1 Diffuser: Diffusion Models for Planning
Diffuser (diffuser) addresses long‐horizon decision-making by treating entire trajectories as a matrix
@@
Accordingly, at each denoising step, gradient information from nudges the model toward trajectories that appear both feasible (as learned from the offline data) and promising with respect to returns, in a manner akin to classifier guidance in image diffusion (classifier_guided).
Diffusion Forcing (diffusion_forcing) extends the above framework by allowing tokenization of . This tokenization allows each token to be denoised at a different noise level, enabling partial denoising of segments where uncertainty is high (e.g., future plans), without requiring a complete transition from full noise to no noise across the entire trajectory. Such token-level control is particularly beneficial in domains that demand causal consistency, such as long-horizon planning problems.
@@ 2.2 Monte Carlo Tree Search
Monte Carlo Tree Search (MCTS) is a planning algorithm that combines tree search with stochastic simulations to effectively balance exploration and exploitation (mcts). Typically, MCTS proceeds in four stages: selection, expansion, simulation, and backpropagation. During selection, the algorithm starts at a root node corresponding to the current state (or partial plan) and descends the tree according to a policy such as Upper Confidence Bounds for Trees (UCT) (uct), aiming to balancely choose the most promising child at each step or explore unexperienced states. Once a leaf or expandable node is reached, expansion adds new child nodes representing previously unexplored actions or plans.
@@ 3.1 Denoising as Tree-Rollout
To address this issue, we first introduce the Denoising as Tree-Rollout process by leveraging the semi-autoregressive denoising process (diffusion_forcing). Specifically, we partition the full trajectory into subplans, such that . Unlike the standard diffuser, where all subplans share the same global denoising schedule, our approach allows assigning independent denoising schedules to each subplan. By applying faster denoising to earlier subplans and slower denoising to later subplans, the process effectively becomes causal and semi-autoregressive. This allows the denoising process to determine the future conditioned on the already determined past. Consequently, while preserving the globally consistent and holistic generation advantages of Diffuser, the denoising process approximates:
Notably, although this formulation appears autoregressive, it is still executed as a single denoising process by controlling noise levels across subplans (diffusion_forcing).
@@ 3.2 Guidance Levels as Meta-Actions
For simplicity, consider two guidance levels: guide and no_guide. In MCTD, we observe that sampling from the prior distribution , i.e., using the standard diffusion sampler represents exploratory behavior as it does not attempt to achieve any goal but only imitates the prior behavior contained in the offline data. Thus, we correspond it to the no_guide meta-action. Conversely, sampling from a goal-directed distribution , such as through classifier-guided diffusion (classifier_guided), represents exploitative behavior, and thus meta-action, guide, is assigned. This steers the sampling process toward achieving a specific goal defined by a reward function .
@@ 3.3 Jumpy Denoising as Fast Simulation
In MCTD, we implement this simulation functionality using the fast jumpy denoising process based on DDIM (ddim). Specifically, when the tree-rollout denoising process has progressed up to the -th subplan, the remaining steps are denoised quickly by skipping every step:
@@ 4 Related Works
Diffusion models (diffusion) have recently shown significant promise for long‐horizon trajectory planning, particularly in settings with sparse rewards, by learning to generate entire sequences rather than stepwise actions (diffuser; diffusion_forcing; decisiondiffuer; d-mpc; hd; plandq; hmd). Various enhancements have been proposed to extend their capabilities. For instance, diffusion_forcing(diffusion_forcing) incorporate causal noise scheduling for semi‐autoregressive planning, and other works introduce hierarchical structures (hd; plandq; hmd), low‐level value learning policies (plandq), or classifier‐free guidance (decisiondiffuer; d-mpc). Despite these developments, the explicit interplay between exploration and exploitation in diffusion sampling has received relatively little attention.
diffusion_forcing further propose Monte Carlo Guidance (MCG) to leverage multiple sub‐plans and average their guidance signals, thereby biasing denoising toward higher‐reward outcomes. While this can encourage planning toward an expected reward over multiple rollouts, it does not implement an explicit search mechanism. The detailed discussion is in Appendix C. Similarly, anonymous2024implicit apply a discrete diffusion denoising process to tasks such as chess, implicitly modeling MCTS without explicit tree expansion.
MCTS (mcts) has achieved impressive results across various decision-making problems, particularly when combined with learned policies or value networks (mcts_alphago; muzero). It has also been applied to System 2 reasoning in Large Language Models (LLMs) to enhance structured problem-solving (xiang2025towards; ToT; zhang2024rest). However, to the best of our knowledge, this work is the first to integrate tree search with diffusion models for full trajectory generation, bridging structured search with generative planning.
@@ 5 Experiments
We evaluate the proposed approach, MCTD, on a suite of tasks from the Offline Goal‐conditioned RL Benchmark (OGBench) (4_park2024ogbench), which spans diverse domains such as maze navigation with multiple robot morphologies (e.g., point‐mass or ant) and robot‐arm manipulation. Our chosen tasks—point and antmaze navigation, multi‐cube manipulation, and a newly introduced visual pointmaze—jointly assess a planner’s ability to handle long‐horizon planning, sequential manipulation, and partial visual observability. In the visual pointmaze, an agent perceives RGB image observations of the 3D environment, thereby testing each method’s resilience to partial observability and the ability to handle image-based planning.
@@ 5.1 Baselines
We compare MCTD against several baselines, each employing a distinct strategy for test‐time computation. First, we include the standard single‐shot Diffuser (diffuser), which plans only at the start of an episode and executes that plan without further adjustments. Next, we consider two methods that allocate additional test‐time compute differently. Diffuser‐Replanning periodically replans at fixed intervals, allowing partial corrections during the episode. This enables us to gauge the benefits of iterative replanning without a full tree search. Diffuser‐Random Search (d-mpc) generates multiple random trajectories in parallel and selects the highest‐scoring one according to the same reward used by MCTD. While this increases test‐time sampling and applies multiple guidance levels for different samples, it lacks the systematic expansion and pruning of a tree‐search algorithm. We also compare our method to Diffusion Forcing (diffusion_forcing), which introduces a causal denoising schedule that enables semi‐autoregressive trajectory generation. This comparison helps isolate the benefits of semi‐autoregressive denoising from our full tree‐structured approach. Further implementation details about all baselines can be found in Appendix A.1.
@@ 5.2 Maze Navigation with Point‐Mass and Ant Robots
We begin with the pointmaze and antmaze tasks from OGBench, featuring mazes of varying sizes (medium, large, and giant). The agent is rewarded for reaching a designated goal region, and the relevant dataset (navigate) comprises long‐horizon trajectories that challenge each method’s capacity for exploration. As in prior work (diffuser), we employ a heuristic controller for the pointmaze environment, whereas the antmaze environment uses a value‐learning policy (plandq) for low‐level control. Appendix A.5 provides more details on this task evaluation.
@@ Results.
In the antmaze domain, MCTD continues to perform nearly perfectly, while baseline performance generally degrades more significantly than in pointmaze, except for the medium-sized settings under Diffusion Forcing. A principal reason for the baseline’s degradation is that the value learning policy (dql) struggles in high‐dimensional control tasks. Even the performance degradations exist due to inaccurate execution, MCTD consistently outperforms baselines and the gap is larger as the maze size is larger.
@@ 5.3 Robot Arm Cube Manipulation
For multi‐cube manipulation tasks from OGBench, a robot arm must move one to four cubes to specific table locations as shown in Figure 4(a). Increasing the number of cubes grows both the planning horizon and the combinatorial complexity. While the planner (MCTD or a baseline) issues high‐level positions for the cubes, a value-learning policy (dql) executes local actions. We augment MCTD with object‐wise guidance by selecting which cube is manipulated at each step, avoiding simultaneous movements of multiple cubes. We note that the object-wise guidance is particularly suited to MCTD because the tree-expansion mechanism naturally supports discrete ”meta-actions” at each node, such as selecting which object to move next. Single-pass diffusion methods, by contrast, attempt to produce a single, holistic trajectory in one go, leaving no opportunity to decide object-by-object. This guidance and additional task‐specific reward shaping are described in Appendix A.6.
@@ Results.
Table 2 shows success rates for one to four cubes. All methods perform comparably on single‐cube tasks, but performance drops significantly for multi‐cube scenarios, especially for standard diffusion baselines that struggle to decide cube ordering and handle subtasks such as stacking. However, MCTD-Replanning shows significant performance gaps from other models for multi-cube scenarios. MCTD exhibits moderate gains (e.g., 22% on two cubes), though it suffers from holistic plan entanglements when multiple objects are involved as shown in Figure 4(b). To address this, MCTD‐Replanning periodically replans, effectively separating each cube’s movements as shown in Figure 4(c). This boosts success on the double‐cube problem from 22% to 50%, and also benefits tasks with more cubes. Another challenge arises from the value-learning policy, which generalizes poorly as the number of cubes increases. Even if MCTD generates a suitable global plan, flawed local control undermines the final outcome. Interestingly, training the policy solely on single‐cube data still allows MCTD‐Replanning to solve a portion (24%) of quadruple‐cube scenarios, illustrating the potential of object‐wise guidance and iterative replanning.
@@ 5.4 Visual Pointmaze
We pre‐encode observations with a Variational Autoencoder (VAE) (vae) to generate latent states suitable for diffusion‐based planning. Because Diffuser’s default heuristic controller is inapplicable to purely visual inputs, we design an inverse dynamics model that maps consecutive latent observations to actions, . For guidance, we employ a position estimator trained on ground‐truth coordinates; however, this estimator does not update the underlying VAE, ensuring that the planner remains in a partially observable regime. The details for this task setting are discussed in Appendix A.7.
@@ A.1 Baselines
Diffuser (diffuser): We use the standard single‐shot Diffuser as our primary diffusion baseline. It generates a full trajectory by denoising from maximum to zero noise in one pass, guided by a return‐predicting function . While it can produce coherent plans, it lacks any mechanism to adapt or refine these plans at test time.
@@
Diffuser-Random Search: We consider a random‐search variant of Diffuser that generates multiple trajectories from the randomly noised plans and picks the best candidate according to a heuristic or learned value function. This approach can be seen as a “Sample-Score-Rank” method proposed in (d-mpc), increasing test‐time compute by drawing more samples, but lacking the systematic exploration or adaptation found in tree‐search paradigms.
Diffusion Forcing: Diffusion Forcing (diffusion_forcing) extends Diffuser by introducing a tokenized, causal noise schedule. Subplans of the trajectory can be denoised at different rates, allowing partial re‐sampling of specific time segments without re‐optimizing the entire plan. In our experiments, we evaluate Diffusion Forcing with the Transformer backbone provided in their official repository, which performed better in our preliminary tests. This method helps isolate the benefit of tokenized partial denoising from our tree‐structured approach. We note that Diffusion Forcing baselines are basically designed to do replanning, so we follow the setting in this evaluation.